Mathematics 242 – Principles of Analysis Definitions Quiz 5 – March 18, 2011

Your Name:

A) (10) What is a subsequence of a sequence {x_{n}}?

B) (10) What is the definition of the meaning of the statement lim_{x→c} f(x) = L?